OpenStudy (anonymous):

What is the slope of the line 2x + 12y = 9?

OpenStudy (anonymous):

What is the slope of the line that passes through (-1,-5) and (3,-3)?

OpenStudy (anonymous):

Does the line 4y + 3x = -8 pass through the point (3, 2)

OpenStudy (anonymous):

x = -9 is the equation of a line whose slope is undefined. (True or False)

OpenStudy (anonymous):

y = -5x + 7 is the equation of a line that passes through the point (3,6) and has a slope of -5. (True or False)

OpenStudy (anonymous):

for the 1st question set the equation in terms of y and you get \[12y=9-2x \rightarrow y=\frac{ 9 }{ 12 }-\frac{ 1 }{ 6 }x\]

OpenStudy (anonymous):

you know what the slope is in the equation right?

OpenStudy (texaschic101):

1. 2x + 12y = 9 write the equation in slope intercept form (y = mx + b)-- m = the slope 12y = -2x + 9 (divide both sides by 12 to get y by itself) 12y/12 = -2/12x + 9/12 y = -1/6x + 3/4 slope of this line is : - 1/6 2. set 1 (-1,-5)...set 2 (3,3) m = (y2 - y1)/(x2 - x1) now sub... m = (3 - (-5)/3 - (-1) m = 8/4 = 2 slope of these points is 2 3. 4y + 3x = - 8 ....points given (3,2) find the slope... y = mx + b 4y = -3x - 8 y = - 3/4x - 2 now use point slope formula y - y1 = m(x - x1) y - 2 = -3/4(x - 3) y - 2 = - 3/4x + 9/4 y = -3/4x + 9/4 + 2 y = - 3/4x + 17/4 no, they do not cross, however they are parallel. 4. x = - 9 IS the equation of a line whose slope is undefined. This line is vertical. 5. points (3,6)....slope -5 y - y1 = m(x - x1) y - 6 = -5(x - 3) y - 6 = -5x + 15 y = -5x + 15 + 6 y = - 5x + 21 False. These lines are parallel.